OpenStudy (anonymous):

Log base 6 of (x+1) = 2? Solve for x.

ganeshie8 (ganeshie8):

are you trying to solve \(x\) ?

OpenStudy (anonymous):

Solving for x

ganeshie8 (ganeshie8):

Okay, do you know how to change a logarithm into exponent form ?

ganeshie8 (ganeshie8):

If \(\log_a b = c \), then \(b = a^c\)

ganeshie8 (ganeshie8):

If \(\log_6 (x+1) = 2 \), then \(x+1 = 6^2\)

ganeshie8 (ganeshie8):

see if you can solve \(x\)

OpenStudy (anonymous):

X=35

ganeshie8 (ganeshie8):

Looks good!

OpenStudy (anonymous):

Thank you! You really helped!

ganeshie8 (ganeshie8):

yw:)